A ball is shot from the top of a building with an initial velocity of 19 m/s at an angle θ = 45 ∘ above the horizontal. Part A What are the x and y components of the initial velocity? Express your answers using two significant figures. Enter your answers numerically separated by a comma. Part B If a nearby building is the same height and 58 m away, how far below the top of the building will the ball strike the nearby building? Express your answer using two significant figures.
Answer #1

vx = 19 cos 45 = 13 m/s

vy = 19 sin 45 = 13 m/s

========

mition along horizontal

d = vx * t

t = 58 / 13.435 = 4.317 s

motion along vertical

- h = vy * t - 0.5 g t^2

-h = - 33.32 m

h = 33 m

=======

Comment in case any doubt, will reply for sure.. Goodluck